OpenStudy (anonymous):

Does x2 - 18x + 45 and x2 - 6x + 9 share a common factor? Please explain too ;~;

OpenStudy (anonymous):

have you tried getting the factors of each of the polynomials?

OpenStudy (anonymous):

I factored them both and got (x - 3)(x - 15) and (x-3)^2 if that's what you were asking.

OpenStudy (solomonzelman):

yes, so if \(\large\color{midnightblue}{ \rm x^2 - 18x + 45 =(x-15)(x-3) }\) and, \(\large\color{midnightblue}{ \rm x@2 - 6x + 9 =(x-3)(x-3) }\) then the factor the BOTH have in COMMON is ?

OpenStudy (anonymous):

3?

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm (x-3) }\)

OpenStudy (anonymous):

oooh okay thank you!

OpenStudy (solomonzelman):

Anytime